PLEASE HELP! :)


How much water must be added to 5 litres of a 90% juice drink to make a 75% juice drink?


i) Calculate the amount of juice and the amount of water in the 5 litres of the 90% juice drink.


ii) Assume x litres of water was added to make a 75% juice drink. Write the algebraic equation for the ratio of juice to total water which gives the 75% juice drink.


iii) Solve the equation for x.


ANY help would greatly be appreciated and welcomed. But PLEASE remember to explain each step thoroughly!


To anyone who attempts this, thank you so much for the help!

Found 2 solutions by josgarithmetic, vindeep:
Answer by josgarithmetic(39621)

Two Part Mixture when One Part Unknown
http://www.algebra.com/my/Two-Part-Mixture-with-one-material-quantity-unknown.lesson?content_action=show_dev

Answer by vindeep(4)

(i) the amount of juice = 90% x 5litre = 4.5 Litre
the amount of water = 10% x 5 litre = 0.5 litre
(ii)amount of water added = x litre
so total quantity of drink = (5+x)litres
ratio of juice to total water = 4.5 / (x+0.5)

4.5 / ( 5 + x) = 75%
(iii) 4.5 = (5+x) 3 /4
4.5 X 4 = 15 + 3x
solving this gives x = 1 litre